14. The number of absences from June to March of a student based on

his class adviser's record is shown below.

Month No. of Absences Month No. of Absences

June

November

2

July

2

December

1

August

1

January

4

September

3

February

1

October

March


If X be the random variable representing the number of absences. Which

table represents the probability distribution?

A.


1

2

3

P(X)

2/10

2/10

3/10

3/10

В.

1

2

3/10

3

3/10

PX)

1/10

3/10

C.

1

3

4

X

P(X)

O

1/10

2

3/10

1/10

3/10

1/5

X

O

1

2

3

PX)

1/5

3/10

3/ 10

1/10

1/10


1
Expert's answer
2022-03-25T06:00:41-0400

"Numbers\,\,of\,\,Absences:\\2,2,2,1,1,4,3,1,0,0\\P\left( X=0 \right) =\frac{2}{10}=\frac{1}{5}\\P\left( X=1 \right) =\frac{3}{10}\\P\left( X=2 \right) =\frac{3}{10}\\P\left( X=3 \right) =\frac{1}{10}\\P\left( X=4 \right) =\frac{1}{10}"

The last table.